OpenStudy (anonymous):

find x such that the distance between (x, -10) and (-20,-5) is the square root of 425.

OpenStudy (anonymous):

\[distance=\sqrt{(x-x1)^2+(y-y1)^2}\]

myininaya (myininaya):

\[\sqrt{(x-(-20))^2+(-10-(-5))}=\sqrt{425}\]

OpenStudy (anonymous):

425=(x+20)^2+(-5)^2

OpenStudy (anonymous):

wait; where did you loose the squared on the (-10+5)

myininaya (myininaya):

fixed lol

myininaya (myininaya):

good eye

OpenStudy (anonymous):

lol it confused me for a sec!

myininaya (myininaya):

\[425=(x+20)^2+(-5)^2\]

myininaya (myininaya):

it was a type-o

myininaya (myininaya):

the 2 should have always been there

OpenStudy (anonymous):

haha okay; i was just making sure!

myininaya (myininaya):

until this step \[425=(x+20)^2+25\]

myininaya (myininaya):

\[400=(x+20)^2\]

myininaya (myininaya):

you will get two solutions x=-20 or x=0 lol i think i should retire for tonight

OpenStudy (anonymous):

haha; thank you so much!